# Context-Aware Multimodal Representation Learning for Spatio-Temporally Explicit Environmental modelling

### v1

@@ Abstract
Earth observation (EO) foundation models have emerged as an effective approach to derive latent representations of the Earth system from various remote sensing sensors. These models produce embeddings that can be used as analysis-ready datasets, enabling the modelling of ecosystem dynamics without extensive sensor-specific preprocessing. However, existing models typically operate at fixed spatial or temporal scales, limiting their use for ecological analyses that require both fine spatial detail and high temporal fidelity. To overcome these limitations, we propose a representation learning framework that integrates different EO modalities into a unified feature space at high spatio-temporal resolution. We introduce the framework using Sentinel-1 and Sentinel-2 data as representative modalities. Our approach produces a latent space at native 10 m resolution and the temporal frequency of cloud-free Sentinel-2 acquisitions. Each sensor is first modeled independently to capture its sensor-specific characteristics. Their representations are then combined into a shared model. This two-stage design enables modality-specific optimisation and easy extension to new sensors, retaining pretrained encoders while retraining only fusion layers. This enables the model to capture complementary remote sensing data and to preserve coherence across space and time. Qualitative analyses reveal that the learned embeddings exhibit high spatial and semantic consistency across heterogeneous landscapes. Quantitative evaluation in modelling Gross Primary Production reveals that they encode ecologically meaningful patterns and retain sufficient temporal fidelity to support fine-scale analyses. Overall, the proposed framework provides a flexible, analysis-ready representation learning approach for environmental applications requiring diverse spatial and temporal resolutions.
@@ I Introduction
EO foundation models have recently emerged as a promising solution [longepe2025earth, galaz2025ai]. Overall, these models aim to learn generalized representations of the Earth system across multiple sensors, regions, and time periods that can be directly applied to diverse downstream tasks. They produce transferable, analysis-ready embeddings, bypassing the need for expert-driven preprocessing and harmonization across sensors. Such embeddings have demonstrated the potential to capture ecosystem dynamics at scale for various modalities, e.g., [fibaek2024phileo, allen2024m3leo]. However, they typically operate at fixed spatio-temporal granularities that limit analytical flexibility. Google’s AlphaEarth [brown2025alphaearth] embeddings provide consistent large-scale representations but operate on fixed, yearly intervals that do not resolve intra-annual variability that is essential for downstream applications such as crop monitoring [murakami2025within, sisheber2024influence] or vegetation phenology [zhang2009sensitivity]. Similarly, TESSERA [feng2025tessera] embeddings compress Sentinel-1 and Sentinel-2 time series into a single annual representation per pixel, offering per-pixel coverage but limited temporal flexibility. Conversely, temporally resolved models such as the Copernicus Foundation Model [wang2025towards] and Major TOM [czerkawski2024global] operate at coarse spatial resolutions (26 km and 10 km, respectively), limiting their utility for fine-grained biodiversity assessments [robertson2023effects], or habitat mapping [cruz2024mapping] where local heterogeneity is crucial. While these embeddings demonstrate high scalability, their spatial or temporal resolution limits their ability to resolve local short-term vegetation dynamics or to adapt to task-specific time horizons. Consequently, their applicability for ecological analyses remains restricted.
To address these limitations, we introduce a multimodal, context-aware representation learning framework that can integrate different EOs into a unified latent feature space.We illustrate our approach using Sentinel-1 radar and Sentinel-2 optical observations. This framework facilitates a complementary structural and spectral encoding while preserving coherence across space and time. By maintaining the native temporal and 10-m spatial resolution of Sentinel-2 observations, our model captures fine-scale ecosystem dynamics. The result is a set of compact, spatio-temporally explicit descriptors of land-surface conditions that can be flexibly applied to diverse Earth system applications.
@@ II Data
To ensure a diverse representation of European ecosystems, we randomly selected 250 training cube locations of $900\times 900~\text{m}$ across different land cover classes using the European Space Agency (ESA) Climate Change Initiative (CCI) land cover map [esa_landcover_2024]. The sampling domain was restricted to Central and Western Europe, specifically the bounding box $0^{\circ},42^{\circ},30^{\circ},62^{\circ}]$ (west, south, east, north). The distribution of selected land cover types is summarized in Table I. We maintained a minimum spatial separation of 50 km among cube centers to reduce spatial autocorrelation.
At each selected location, we extracted two annual Sentinel-2 Level-2A [drusch2012sentinel] data cubes from acquisition periods separated by at least one year. Each of the resulting 500 cubes comprises 12 spectral bands (B01, B02, B03, B04, B05, B06, B07, B08, B8A, B09, B11, B12). For model training, only the ten Sentinel-2 bands at 10 and 20 m spatial resolution were retained, excluding the 60 m bands (B01 and B09). Each cube covers a continuous 12-month period between 2017 and 2024. This temporal block sampling reduces temporal biases and ensures coverage across different seasonal and inter-annual conditions. The temporal resolution follows the Sentinel-2 revisit cycle, yielding approximately 130–150 observations per year in Central Europe. To prevent data leakage, all samples derived from a single cube were used exclusively for either training, validation, or testing.
@@
We retrieved the corresponding spatially and temporally aligned Sentinel-1 data from Microsoft Planetary Computer [^1] https://planetarycomputer.microsoft.com/dataset/{Sentinel-1}-rtc. This collection provides radiometrically terrain-corrected (RTC) radar data [small2011flattening], which uses PlanetDEM [^2] https://www.planetobserver.com/global-elevation-data as the elevation reference for normalization. To further reduce speckle noise and short-term fluctuations caused by rainfall or varying sensor geometry, we applied a three-frame rolling mean to the RTC time series in the linear scale.
The full dataset (composed of date from both sensors) was divided into a training set (375 cubes, 75%), a validation set (83 cubes, 17%), and a test set (42 cubes, 8%) to ensure a reliable model evaluation.
@@ III Methods
During this staged approach, we use a context-aware learning strategy, in which the reconstruction objective emphasizes the central coordinate while simultaneously incorporating information from its spatial and temporal neighborhood. This allows the network to exploit the dependencies among nearby observations, resulting in representations that are aligned with real-world structures.
@@ III-A Context-Aware Learning
Our model adopts a context-aware training strategy that emphasizes the central spatio-temporal coordinate and softly incorporates its neighborhood (Fig. 1). To achieve this, the autoencoder processes spatio-temporal patches of $15\times 15$ pixels across 11 time steps. The central pixel serves as the reconstruction target, while its surroundings provide contextual cues. Observations that are spatially or temporally closer receive higher weighting during training, reflecting their stronger ecological relevance. This encourages the model to produce features that are locally specific yet contextually informed, aiming to improve robustness to noise and spatio-temporal irregularities.
@@
where $\mu_{x}$ and $\mu_{\hat{x}}$ are the local means, $\sigma_{x}$ and $\sigma_{\hat{x}}$ the variances, and $\sigma_{x\hat{x}}$ the covariance between the original and reconstructed patches. Constants $c_{1}$ and $c_{2}$ stabilize the division in regions of low variance. In the loss formulation, SSIM is converted into a distance measure:
@@ III-C Data Fusion
During modality fusion, the pretrained encoders and decoders remained frozen, while only the newly introduced fusion layers were optimized. Latent features from the Sentinel-1 and Sentinel-2 encoders were first passed through modality-specific linear projection layers and then concatenated. The combined features were processed by a Transformer encoder equipped with temporal positional encodings, enabling the model to account for asynchronous radar–optical acquisitions and to capture cross-modal temporal dependencies. Subsequently, a linear projection layer produced a compact seven-dimensional bottleneck representation encoding the central pixel within a temporally consistent, multimodal embedding space. These fused representations were then projected back through linear layers and reconstructed by the respective decoders.
@@ IV-A1 Modality Pretraining Results
Quantitatively, the weighted MAE for Sentinel-1 converged to 0.0009 (training) and 0.0006 (validation) for the central pixel, and to 0.0089 (training) and 0.0319 (validation) for the surrounding context. For Sentinel-2, the central-pixel MAE reached 0.0037 (training) and 0.0037 (validation), while SAM converged to 0.0259 rad ($\approx$1.5) in training and 0.0278 rad ($\approx$1.6) in validation, confirming minimal angular deviation between reconstructed and reference spectra. The SSIM loss stabilized at 0.1476 (training) and 0.1431 (validation), and the contextual weighted MAE reached 0.0064 (training) and 0.0141 (validation). Overall, the close correspondence between training and validation metrics across all loss components highlights the generalization capability of the autoencoders, reconstructing the target pixel while embedding surrounding context within each modality.
@@ IV-C GPP Modelling
Montero et al. [montero2024recurrent] reported a validation NRMSE of about 0.13 when predicting GPP across 19 forest sites using multimodal observations. While our results show lower errors, the setups are not directly comparable, as we train on fewer sites and slightly different vegetation types. Another realated study [pabon2022potential] estimated based on Sentinel-2 spectral indices and bands GPP from 58 EC sites. Their machine-learning approach achieved RMSE values between 2.5 and 3.0. These results indicate that our learned multimodal embeddings effectively capture relevant information for predicting carbon fluxes, providing performance comparable to models to previous approaches
@@ V Discussion
The resulting representations enable flexible analyses for a broad range of downstream applications that require different spatial and temporal resolutions, from short-term vegetation dynamics to long-term ecosystem trends. Because the embeddings retain explicit spatio-temporal structure at high spatial resolution (10m) and follow the natural acquisition frequency of the sensors after cloud masking, they can be aggregated or sampled flexibly to match diverse ecological and climate-modelling requirements.
In our context-aware loss function, we chose $\alpha=0.1$ to prioritise the immediate neighborhood while still preserving broader contextual cues. Although this choice yielded stable training and consistent results across modalities, we did not experiment with different $\alpha$ values, leaving a systematic exploration of its sensitivity for future work. More generally, other hyperparameters, such as the Sentinel-2 reconstruction loss weights during pretraining or the fusion loss configuration were determined through limited empirical experimentation rather than extensive tuning. We did not perform a comprehensive hyperparameter optimization due to hardware limitations, as we used only a single GPU for training. Nevertheless, the framework demonstrated stable convergence and reliable reconstruction across modalities.
@@
The good agreement between the fused embeddings and GPP dynamics indicates that the proposed model has learned physically meaningful representations of the land surface. The Sentinel-2 component is primarily sensitive to canopy chlorophyll and water status through its near-infrared (NIR) and short-wave infrared (SWIR) bands, which have long been used to infer vegetation density and leaf water content [Ceccato et al., 2001; Jackson et al., 2004]. In contrast, the Sentinel-1 backscatter carries complementary information on vegetation structure, biomass, and surface moisture due to the sensitivity of C-band microwave signals to dielectric and geometric properties [Dubois et al., 1995; Frison et al., 2018]. By integrating these complementary modalities within a context-aware temporal framework, the model likely encodes key biophysical processes—such as canopy moisture availability, structural complexity, and seasonal phenology—that jointly regulate photosynthetic productivity. The improved GPP prediction accuracy therefore arises not merely from statistical fitting but from the model’s capacity to represent underlying ecological and physical processes.
Unlike fixed-scale products that provide a single yearly representation per pixel or embeddings at continental spatial scales, our high-resolution features maintain high temporal and spatial flexibility. This allows users to adapt the representations to their needs, whether by aggregating over larger areas or selecting varying temporal windows. Such flexibility enables analyses that capture local short-term processes, such as the 90-day sliding-window approach used for GPP modelling in this study, while still supporting regional-to-continental ecological assessments. The model reproduced site-specific productivity patterns with performance comparable to previous multispectral and multimodal GPP studies [pabon2022potential, montero2024recurrent].
In contrast, yearly embeddings would require predicting daily GPP values for the entire year simultaneously. They would also provide too few samples to effectively learn temporal relationships when using our selected sites and time periods. Additionally, previous GPP studies at moderate spatial resolutions remained limited in capturing fine-scale ecological variability. [badgley2019terrestrial, camps2021unified, wang2021tracking], Hence, sufficiently detailed spatial and temporal resolutions are relevant for capturing vegetation productivity. This ability is directly supported by our approach.
A current limitation is that pixel-level embeddings occasionally contain temporal and spatial gaps, as missing acquisitions from Sentinel-1 or cloud-filtered Sentinel-2 propagate into the feature space. In some periods, especially during winter or persistent cloud cover, these gaps can span days to weeks. Addressing this challenge remains an open research avenue. A potential solution is the development of deep learning–based spatio-temporal gap-filling frameworks that can reconstruct missing observations consistently across space and time [weiss2014effective, wang2022remote]. Alternatively, incorporating additional observation sources that not only fill information gaps but also enable learning under incomplete multimodal conditions, where missing data from one sensor can be compensated by complementary information from others [chen2024novel].
@@
Integrating additional sensors, however, introduces further complexity due to differences in spatial resolution [schowengerdt2006remote]. Variations in ground sampling and geometric accuracy can cause scale-dependent inconsistencies and spatial artifacts when fusing heterogeneous data [guo2024obsum, wang2021blocks]. These effects may reduce the spatial coherence of the learned representations, particularly in heterogeneous landscapes. Scale-aware processing strategies, such as adaptive resampling [arun2014intelligent], or attention-based fusion mechanisms [gao2023cross], can help mitigate these issues and ensure consistent integration across resolutions.
@@
Overall, our framework contributes significantly to the development of general-purpose representations for environmental modelling. Both the qualitative analysis and the predictive experiments confirm that the learned Sentinel-1 and Sentinel-2 embeddings encode ecologically meaningful information, capturing spatial structure, surface characteristics, and seasonal vegetation dynamics across diverse landscapes. This demonstrates that the learned feature space provides a transferable, flexible, and scalable foundation for data-driven Earth system science, capable of supporting a wide range of downstream applications that require various spatial resolutions and meaningful temporal context.

### v2

@@ Abstract
Earth observation (EO) foundation models have emerged as an effective approach to derive latent representations of the Earth system from various remote sensing sensors. These models produce embeddings that can be used as analysis-ready datasets, enabling the modelling of ecosystem dynamics without extensive sensor-specific preprocessing. However, existing models typically operate at large spatial or temporal contexts, limiting their use for ecological analyses that require both fine spatial detail and high temporal fidelity. To overcome these limitations, we propose a representation learning framework that integrates different EO modalities into a unified feature space at high spatio-temporal resolution. We introduce the framework using Sentinel-1 and Sentinel-2 data as representative modalities. Our approach produces a latent space at native 10 m resolution and the temporal frequency of cloud-free Sentinel-2 acquisitions. Each sensor is first modeled independently to capture its sensor-specific characteristics. Their representations are then combined into a shared model. This two-stage design enables modality-specific optimisation and easy extension to new sensors, retaining pretrained encoders while retraining only fusion layers. This enables the model to capture complementary remote sensing data and to preserve coherence across space and time. Qualitative analyses reveal that the learned embeddings exhibit high spatial and semantic consistency across heterogeneous landscapes. Quantitative evaluation in modelling Gross Primary Production reveals that they encode ecologically meaningful patterns and retain sufficient temporal fidelity to support fine-scale analyses. Overall, the proposed framework provides a flexible, analysis-ready representation learning approach for environmental applications requiring diverse spatial and temporal resolutions.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
While these embeddings demonstrate high scalability, their spatial or temporal design constraints limit their ability to resolve local, short-term vegetation dynamics or to adapt to task-specific time horizons. Consequently, their applicability for ecological analyses remains restricted.
EO foundation models have recently emerged as a promising solution [longepe2025earth, galaz2025ai]. Overall, these models aim to learn generalized representations of the Earth system across multiple sensors, regions, and time periods that can be directly applied to diverse downstream tasks. They produce transferable, analysis-ready embeddings, bypassing the need for expert-driven preprocessing and harmonization across sensors. Such embeddings have demonstrated the potential to capture ecosystem dynamics at scale for various modalities, e.g., [fibaek2024phileo, allen2024m3leo]. However, they typically operate at fixed spatio-temporal granularities that limit analytical flexibility. Google’s AlphaEarth [brown2025alphaearth] embeddings provide consistent large-scale representations but operate on yearly intervals that do not resolve the intra-annual variability essential for downstream applications, such as analysing fast-changing vegetation dynamics [wen2025detection] or capturing phenological processes [simmonds2019cue, shutt2019environmental]. Similarly, TESSERA [feng2025tessera] embeddings compress Sentinel-1 and Sentinel-2 time series into a single annual representation per pixel, offering per-pixel coverage but limited temporal flexibility. Conversely, temporally resolved approaches such as the Copernicus Foundation Model [wang2025towards] and Major TOM [czerkawski2024global] operate at large spatial contexts ($26\times 26~\text{km}^{2}$ and $10\times 10~\text{km}^{2}$, respectively), limiting their utility for fine-grained biodiversity assessments [robertson2023effects], or habitat mapping [gregovich2025vegetation, de2021identifying] where local heterogeneity is crucial. While these embeddings demonstrate high scalability, their spatial and temporal constraints limit their ability to capture local, short-term vegetation dynamics or to match the specific spatial or temporal horizons required by ecological tasks. Consequently, their applicability for ecological analyses remains restricted.
To address these limitations, we propose a multimodal, context-aware representation learning framework that integrates heterogeneous EO modalities into a unified latent feature space at high spatio-temporal resolution. In this work, we demonstrate the framework using Sentinel-1 radar and Sentinel-2 optical data as representative modalities. The approach enables to encode complementary structural and spectral information while preserving coherence across space and time. By maintaining the native 10 m spatial resolution and the temporal fidelity of cloud-free Sentinel-2 observations, it provides compact, spatio-temporally explicit descriptors of land-surface conditions that can be flexibly applied to diverse Earth system applications.
@@ II Data
To ensure a diverse representation of European ecosystems, we randomly selected 250 training cube locations of $900~\text{m}\times 900~\text{m}$ across different land cover classes using the European Space Agency (ESA) Climate Change Initiative (CCI) land cover map [esa_landcover_2024]. The sampling domain was restricted to Central and Western Europe, specifically the bounding box $0^{\circ},42^{\circ},30^{\circ},62^{\circ}]$ (west, south, east, north). The distribution of selected land cover types is summarized in Table I. We maintained a minimum spatial separation of 50 km among cube centers to reduce spatial autocorrelation.
At each selected location, we extracted two annual Sentinel-2 Level-2A [drusch2012sentinel] data cubes from acquisition periods separated by at least one year. For model training, we use the ten Sentinel-2 spectral bands at 10 and 20 m spatial resolution were retained (B01, B02, B03, B04, B05, B06, B07, B08, B8A, B11, B12). Each cube covers a continuous 12-month period between 2017 and 2024. This temporal block sampling reduces temporal biases and ensures coverage across different seasonal and inter-annual conditions. The temporal resolution follows the Sentinel-2 revisit cycle, yielding approximately 130–150 observations per year in Central Europe.
@@
We retrieved the corresponding spatially and temporally aligned Sentinel-1 data from Microsoft Planetary Computer [^1] https://planetarycomputer.microsoft.com/dataset/Sentinel-1-rtc. This collection provides radiometrically terrain-corrected (RTC) radar data [small2011flattening], which uses PlanetDEM [^2] https://www.planetobserver.com/global-elevation-data as the elevation reference for normalization. To further reduce speckle noise and short-term fluctuations caused by rainfall or varying sensor geometry, we applied a three-frame rolling mean to the RTC time series in the linear scale. Finally, we normalised Sentinel-1 data and Sentinel-2 reflectance to the same 0–1 range.
To prevent data leakage, all samples derived from a single cube were used exclusively for either training, validation, or testing. The full dataset (composed of date from both sensors) was divided into a training set (375 cubes, 75%), a validation set (83 cubes, 17%), and a test set (42 cubes, 8%) to ensure a reliable model evaluation.
@@ III Methods
Learning spatio-temporally explicit EO representations at high resolution requires integrating heterogeneous sensor modalities. Different EO missions capture complementary but fundamentally distinct physical signals, such as radar from Sentinel-1 and multispectral reflectance from Sentinel-2, which vary strongly in dynamic range, noise characteristics, and spatio-temporal sensitivity [conti2025advancing]. Because each modality tends to generalise under different conditions, end-to-end joint training with a single optimisation strategy is often sub-optimal [wang2020makes, sun2021learning, yao2022modality]. Consequently, multimodal learning must reconcile modality-specific structures with cross-modal relationships while maintaining stable optimisation and coherent representations.
@@
During this staged approach, we use a context-aware learning strategy, in which the reconstruction objective emphasizes the central coordinate while simultaneously incorporating information from its spatial and temporal neighborhood. This enables the model to incorporate dependencies among nearby observations, allowing the model to use surrounding information as additional cues.
The following subsections detail the core elements of this framework.
@@ III-A Context-Aware Learning
Our model adopts a context-aware training strategy that emphasizes the central spatio-temporal coordinate and softly incorporates its neighborhood (Fig. 1). To achieve this, the autoencoder processes spatio-temporal patches of $15\times 15$ pixels across 11 time steps. The central pixel serves as the reconstruction target, while its surroundings provide contextual cues. . Observations that are spatially or temporally closer receive higher weighting during training, reflecting their stronger ecological relevance. This encourages the model to produce features that are locally specific yet contextually informed, aiming to improve robustness to noise and spatio-temporal irregularities.
@@
where $\mu_{x}$ and $\mu_{\hat{x}}$ are the local means, $\sigma_{x}$ and $\sigma_{\hat{x}}$ the variances, and $\sigma_{x\hat{x}}$ the covariance between the original and reconstructed patches. Constants $c_{1}$ and $c_{2}$ stabilize the division in regions of low variance. In the loss formulation, SSIM is converted into a discrepancy measure:
@@ III-C Data Fusion
During modality fusion, the pretrained encoders and decoders remained frozen, while only the newly introduced fusion layers were optimized. Latent features from the Sentinel-1 and Sentinel-2 encoders were first passed through modality-specific linear projection layers and then concatenated. The combined features were processed by a Transformer encoder equipped with temporal positional encodings, enabling the model to account for asynchronous radar–optical acquisitions and to capture cross-modal temporal dependencies. Subsequently, a linear projection layer produced a compact seven-dimensional bottleneck representation encoding the central pixel within a temporally consistent, multimodal embedding space. The bottleneck dimensionality was chosen empirically as it achieved the lowest reconstruction MAE for the central pixel of the spatio-temporal patch. These fused representations were then projected back through linear layers and reconstructed by the respective decoders.
@@ IV-A1 Modality Pretraining Results
Quantitatively, the weighted MAE for Sentinel-1 converged to 0.0009 (training) and 0.0006 (validation) for the central pixel, and to 0.0324 (training) and 0.0319 (validation) for the surrounding context. For Sentinel-2, the central-pixel MAE reached 0.0037 (training) and 0.0037 (validation), while SAM converged to 0.0259 rad ($\approx$1.5) in training and 0.0278 rad ($\approx$1.6) in validation, confirming minimal angular deviation between reconstructed and reference spectra. The SSIM loss stabilized at 0.1476 (training) and 0.1431 (validation), and the contextual weighted MAE reached 0.0142 (training) and 0.0141 (validation). Overall, the close correspondence between training and validation metrics across all loss components highlights the generalization capability of the autoencoders, reconstructing the target pixel while embedding surrounding context within each modality.
@@ IV-C GPP Modelling
Montero et al. [montero2024recurrent] reported a validation NRMSE of about 0.13 when predicting GPP across 19 forest sites using multimodal observations. While our results show lower errors, the setups are not directly comparable, as we train on fewer sites and slightly different vegetation types. Another related study [pabon2022potential] estimated GPP based on Sentinel-2 spectral indices and bands from 58 EC sites. Their machine-learning approach achieved RMSE values be tween 2.5 and 3.0. These results indicate that our learned multimodal embeddings effectively capture relevant information for predicting carbon fluxes, providing performance comparable to models to previous approaches.
@@ V Discussion
The resulting representations enable flexible analyses for a broad range of downstream applications that require different spatial and temporal contexts, from short-term vegetation dynamics to long-term ecosystem trends. Because the embeddings retain explicit spatio-temporal structure at 10m resolution and follow the natural acquisition frequency of the sensors after cloud masking, they can be aggregated or sampled flexibly to match diverse ecological and climate-modelling requirements.
In our context-aware loss function, we chose $\alpha=0.1$ to prioritise the immediate neighborhood while still preserving broader contextual cues. Although this choice yielded stable training and consistent results across modalities, we did not experiment with different $\alpha$ values, leaving a systematic exploration of its sensitivity for future work. More generally, other hyperparameters, such as the Sentinel-2 reconstruction loss weights during pretraining or the fusion loss configuration were determined through limited empirical experimentation rather than extensive tuning. We did not perform a comprehensive hyperparameter optimization due to hardware constraints. Nevertheless, the framework demonstrated stable convergence and reliable reconstruction across modalities.
@@
Using the learned multimodal embeddings, the model successfully reproduced site-specific GPP dynamics with accuracy comparable to previous multispectral and multimodal approaches [pabon2022potential, montero2024recurrent]. These results indicate that the representations capture information relevant for modelling photosynthetic activity across diverse vegetation types and environmental conditions. The ability to predict daily GPP consistently across different sites demonstrates that the fused feature space encodes meaningful temporal signals that relate to ecosystem functioning.
The agreement between the fused embeddings and GPP dynamics indicates that the model has learned physically grounded representations of the land surface. Sentinel-2 contributes spectral signals related to canopy chlorophyll, vegetation density, and leaf water content through its near-infrared and short-wave infrared bands [ceccato2001detecting, jackson2004vegetation]. Sentinel-1 complements this with sensitivity to vegetation structure, biomass, and surface moisture owing to the dielectric and geometric response of C-band microwaves [dubois2020characterization, frison2018potential]. Together, these modalities encode biophysical variability closely associated with seasonal productivity patterns across sites.
A major advantage of the learned embeddings is their high temporal and spatial flexibility. Users can aggregate them over larger spatial units or select task-specific temporal windows depending on the application. This flexibility enabled the 90-day sliding-window setup used in our GPP experiment, which captures short-term seasonal dynamics while preserving 10 m spatial detail.
In contrast, yearly embeddings require predicting all daily GPP values simultaneously from a single latent representation. This provides too few samples per site to learn robust temporal relationships. Previous work has shown that the EC footprint varies substantially across seasons and sites. Consequently, approaches that rely on coarse spatial resolution can introduce representativeness biases, reducing the reliability of GPP analyses [schmid1997experimental, chen2009assessing, chu2021representativeness]. High spatial resolution provides the flexibility needed to select footprint-aligned areas, ensuring that the domain used for model assessment closely matches the conditions actually observed by the tower [chu2021representativeness]. Our spatio-temporally explicit approach directly addresses both limitations.
A current limitation is that the learned features contain temporal and spatial gaps, as missing acquisitions from Sentinel-1 or cloud-filtered Sentinel-2 propagate into the feature space. In some periods, especially during winter or persistent cloud cover, these gaps can span days to weeks. Established gap-filling techniques offer practical ways to mitigate such effects. These include spatial interpolation approaches such as kriging [addink1999comparison, rossi1994kriging] or temporal smoothing methods including the Savitzky–Golay and seasonal filters implemented in TIMESAT [jonsson2004timesat], and hybrid spatio-temporal strategies that alternate between spatial and temporal modelling steps [kang2005improving, borak2009effective]. Alternatively, machine learning methods can be applied [weiss2014effective].
@@
To simplify the integration of additional sensors, our approach first models each modality independently and trains only the fusion layers jointly. This design allows every encoder to be optimised with modality-specific learning strategies while keeping the overall framework flexible and extensible. A remaining challenge arises from differences in spatial resolution. Variations in ground sampling and geometric accuracy can introduce scale-dependent inconsistencies and spatial artefacts when combining heterogeneous data [schowengerdt2006remote, guo2024obsum, wang2021blocks]. These effects may reduce spatial coherence, particularly in heterogeneous landscapes. However, scale-aware processing strategies, such as adaptive resampling [arun2014intelligent] or attention-based fusion mechanisms [gao2023cross], provide effective ways to mitigate these issues and support consistent integration across resolutions.
@@
Overall, our framework contributes significantly to the development of general-purpose representations for environmental modelling. Both the qualitative analysis and the predictive experiments confirm that the learned Sentinel-1 and Sentinel-2 embeddings encode ecologically meaningful information, capturing spatial structure, surface characteristics, and seasonal vegetation dynamics across diverse landscapes. This demonstrates that the learned feature space provides a transferable, flexible, and scalable foundation for data-driven Earth system science, capable of supporting a wide range of downstream applications that require various spatial and temporal context of high resolution.